### v1

@@ 3.6 Double peak lines: Jet detection
Double-peaked spectral lines in protostars are linked to complex geometries and kinematics. These lines typically arise from Doppler-shifted emission caused by fast moving flows with their components moving at different velocities. High-resolution observations of many protostars have shown that forbidden emission lines often exhibit two velocity components, namely the low-velocity (LV) and high-velocity component, see review by Eislöffel et al. (2000). Such velocity separations can also arise in bow-shock structures, where a curved shock front produces a range of shock velocities and spatially separated emission regions (Hartigan et al., 1987). These interpretations have been supported by theoretical simulations with magnetohydrodynamic (MHD) jet models (Garcia et al., 2001; Shang et al., 2002). They have shown through their position-velocity diagrams the presence of different velocity components.
@@
Closer inspection of the L1527 MIRI/MRS spectra reveals clear double peak emission in three different forbidden lines: [Ar ii], [Ne iii], and [Fe ii]. The [Ne ii] line showed very weak and un-resolved profile limited to just few pixels and was therefore excluded. The double peak emission was detected only on the eastern region of the outflow. Fig. 8 shows the profiles and the morphology of double peak spectra for [Ar ii], [Ne iii], and [Fe ii] lines. In the first column, the Moment 0 maps are presented with the sample aperture positions. In the middle column, we show the double peak spectra extracted at the sample aperture position. The spectra are color coded blue, green, and red for [Ar ii], [Ne iii], and [Fe ii], respectively. The line profiles clearly show two separate velocity peaks associated to LV and HV components. A 1-D Gaussian fit was performed on each velocity component to estimate their peak radial velocity, shifted to the rest frame in LSR. In the last column of Fig. 8, we show the jet emission map associated to the HV component. The jet emission was isolated from the double-peaked line profile by extracting the high-velocity (HV) component, thereby identifying the spatial region associated with it. This was achieved by integrating the line intensity over a velocity range of km s-1, with an additional criteria that the emission exhibited a Gaussian profile and a line width exceeding the spectral resolution. All the three maps show an extended jet with its brightest emission closest to the protostar.
@@
The LV component observed in the double peak lines is distributed throughout the region and associated to the shocked gas in the general outflow or disk wind. The average radial velocity of the LV component is about km s-1, which is notably blue shifted. However, we can confirm that in regions without double peaked profiles, the average radial velocity is red shifted and is consistent with the [Ne ii] velocity maps. Disk winds launched from few to several au’s from the protostar are expected to have velocities of tens of km s-1(Pudritz et al., 2007; Shang et al., 2007). They are thought to shape the molecular outflow cavities by sweeping up surrounding envelope material. Our observed LV component velocities are consistent with these predictions and supports previous studies indicating that the L1527 outflow cavity is shaped by a slow, disk wind entraining the ambient material (Tobin et al., 2008, 2012).
@@
The double peak HV component is present only in the eastern region of the outflow. The [Ar ii] line shows the strongest emission among the three detected lines. The average radial velocity of the HV component is about km s-1. The emission is not narrow/collimated, rather it is extended in width and follows the direction of the outflow axis to about 500 au. We see that the emission bends/curves towards the south in all the three maps. The HV component is interpreted as tracing the presence of a jet. Although the emission is not strictly collimated as seen in other Class 0/I jets, its observed velocity (100 km s-1), extended scale, and detection in high-ionization lines ([Ar ii], [Ne iii]) strongly suggest shock-excited jet activity, a strong characteristic of early-stage protostellar evolution.
@@
Considering the jet is almost in the plane-of-sky, we can correct the radial velocity for inclination. For a value of , we get the true velocity to be around km s-1. This value should be interpreted cautiously as the disk is warped and the jet inclination would vary significantly due its precession/bending. Very high velocities (200 km s-1) are predicted by magneto-centrifugal launching models when jets originate very close to the protostar. X-wind models (Shu et al., 2007), which launch material near the co-rotation radius, naturally produce such high velocities due to the strong gravitational potential in this region. However, disk-wind models (Ferreira , 1997) can also generate comparably high velocities when launched from sub-au radii. Therefore, the large velocities observed in the jet likely indicates that its launching region lies very close to the protostar. Overall, the coexistence of both velocity components in multiple ionized lines suggests L1527 is composed of a stratified layer of gas with varying excitation conditions.
@@ 4.1 Temperature and Column density maps
The physical properties like gas temperature , and the column density (H2) can be inferred using excitation diagram analysis of H2 rotational transitions. The observed extinction-corrected H2 line integrated intensities are converted into upper state column densities normalized by its statistical weight (/). These are plotted in a logarithm scale against their upper state excitation energies (/). For gas in local thermal equilibrium (LTE), the gas excitation follows a Boltzmann distribution (//) and the data points in the diagram can be fitted with a straight line. The gas excitation temperature is derived as the reciprocal of the slope of the line, whereas the column density is proportional to the y-axis intercept. This analysis assumes that all H2 transitions are optically thin, which is valid even for (H2)1023 cm-2 (Bitner , 2008).
@@
The MIRI/MRS spectral range covers several pure rotational H2 transitions from the S(8) to S(1) state, which are detected in most pixels in the entire field-of-view. This allows us to carry out pixel-by-pixel excitation diagram analysis and to obtain a complete map of , and (H2). We first smooth the spectral cubes from different MIRI channels to the lowest common resolution as the angular resolution varies from 02 to 10 between all the MIRI channels. We smooth the spectral cubes to a resolution of 07 that corresponds to the S(1) transition, which is the lowest angular resolution in the detected H2 lines. We also regrid the cubes to the same spatial grid as the H2 S(1) data.
@@
Next, we correct for extinction along the line-of-sight. The most commonly used extinction correction in MIR comes from McClure (2009) and KP5 (Pontoppidan et al., 2024) curves, with the latter introducing slightly stronger absorption from ices and silicates. Previous analysis of other JOYS sources adopted the McClure (2009) extinction curve that is valid between a -band magnitude of 1 to 7 mag. However, recent study on ices in L1527 by Slavicinska et al. (2025) showed substantial ice and silicate absorption, indicating that the KP5 curve provides a more appropriate extinction correction, which we apply in the analysis. The optical depth at 9.7m was earlier estimated to be around (Slavicinska et al., 2025). We also measured the extinction by de-reddening the S(1) to S(8) lines on excitation diagram analysis, which yielded mag (see Fig. 13). Based on these results, as well as comparing previous extinction derived with SED modeling (Tobin et al., 2008, 2010), we adopt a value of mag, that is roughly equivalent to a visual extinction of mag. We tested lower extinction values between mag and found that the intensity of the S(3) line decreases significantly and gets excluded in the fitting. We also find that an H2 ortho-to-para ratio of 3 provides the best fit to our observed line intensities, assuming the H2 gas is in LTE.
@@
The excitation diagram analysis is performed using the pdrtpy222pdrtpy is developed by Marc Pound and Mark Wolfire. This project is supported by NASA Astrophysics Data Analysis Program grant 80NSSC19K0573; from JWST-ERS program ID 1288 provided through grants from the STScI under NASA contract NAS5-03127; and from the SOFIA C+ Legacy Project through a grant from NASA through award #208378 issued by USRA. python package, which is commonly employed in the analysis of JOYS sources (Gieser et al., 2023; Caratti o Garatti et al., 2024; Francis et al., 2025). Typically, the observed H2 rotational transitions trace two temperature components, often parameterized as a ‘warm’ and a ‘hot’ component in excitation diagrams. The warm component ( K) is traced by the lower-excitation lines (e.g., S(0)-S(4), 4000 K) and is commonly associated with slow, nondissociative -shocks or the cooler post-shock gas within -shocks. The hot component ( K), traced by the higher excitation transitions (e.g., S(5) and above, 5000-10000 K), arises mostly from energetic fast dissociative -shocks where collisional heating dominates. However, recent shock models (e.g., Kristensen et al., 2023) demonstrate that a single shock structure can produce both temperature components, where hot gas represents post-shock region and the warm gas traces the cooling layer of the same shock. Therefore, the presence of two temperature components does not necessarily indicate separate shocks, but can instead reflect a thermally stratified region with distinct excitation conditions. In general, the warm component typically carries most of the H2 luminosity and therefore remains a dominant tracer of the bulk outflowing gas, while the hot component highlights localized regions of strong shock activity.
@@
At the initial step, the extinction corrected H2 transitions at each spaxel are fitted by a 1-D Gaussian and the line intensities are computed from the integral of the Gaussian fit. We set an intensity threshold of 20 MJy sr-1 to avoid any spurious detection. In Fig. B, we show example spectra for spaxels associated to the apertures in the eastern and western lobes (see Fig. 1). The spectra are plotted for different H2 transitions and their Gaussian fits are highlighted in red and blue for eastern and western lobes, respectively. We see that the line intensities vary in value across different transitions depending on their strength. In general, lines from all transitions have bonafide profiles for the Gaussian fits without any omission.
@@
Next, the line intensities from the spectra are converted to upper state column densities and are then plotted against their excitation temperature as described earlier. To have reliable estimates from the excitation diagram, fitting is performed only if five or more transitions are detected in a spaxel. Fig. 14 shows the excitation diagram for the example spaxels in the eastern and western lobes. The data points with error bars are the upper state column densities of lines in each H2 transitions. They are distributed across the plot with overlay of two fits in red and blue comprising of the warm and hot component. The total fit on the data points is shown as a solid green line. The estimated warm temperature in both the spaxels is similar, with an average of about 570 K. Whereas, the hot temperature varies considerably with values of 2170 K and 3340 K for eastern and western spaxels, respectively. The estimated warm and hot column density in both the spaxels average about and .
@@
Fig. 9 presents the full temperature and column density maps for both the warm and hot components. In the temperature maps, the warm component exhibits moderate temperature variations across the outflow region, with values ranging between K. There is a noticeable gradient in the eastern lobe with increasing temperature towards north-east. This gradient closely follows the spatial distribution of the H2 emission, particularly resembling the brighter S(5) and S(4) transitions (see Fig. 3). This suggests that there is presence of warm gas inside the outflow cavity. The western lobe also has a gradient, but with increasing temperature towards south-west and with lesser resemblance to the H2 emission. In the hot component map, the temperature variations are substantial ranging between K. Both sides of the region exhibit sharp increase in temperature along the outflow structures, closely resembling the distribution of the H2 emission. These temperature variations and their spatial resemblance to the outflow morphology suggest that the heating of the gas is strongly linked to a common excitation mechanism within the outflow.
@@
The column density maps exhibit a morphology broadly similar to the temperature maps. The warm component of the column density ranges from , whereas the hot component ranges from . The warm component is significantly denser, about two orders of magnitude higher than the hot component. In the warm component maps, regions of higher column density show a general positive correlation to regions with higher temperature. However, this is in contrast in the hot component maps, where regions of lower column density are seen to present higher temperatures. These trends likely reflect the nature of shock heating in the outflow.
@@ 4.2 Comparison to shock models
The H2 emission from outflows in dense regions ( cm-3) is generally thought to be dominated by collisional excitation from shocks. However, UV driven irradiation may still play a role if accretion-generated UV photons penetrate the base of the outflow cavity before being attenuated, or if fast dissociative shocks generate local UV fields that modify the shock structure and chemistry. The strong gradients in temperature, density, and ionization across shock fronts naturally produce multiple excitation components, with hotter gas populating higher energy rotational states and warmer gas dominating lower levels. In order to constrain the mechanism of excitation in the outflow, we compare our observed H2 emission with recent shock models that self-consistently treat different excitation processes.
@@
We use the recent shock model grid from Kristensen et al. (2023), based on the Paris-Durham code (Flower & Pineau des Fortês , 2003; Godard et al., 2019), which computes the dynamical, thermal, and chemical structure of steady, plane-parallel shocks over 14,000 models. The code determines pre-shock equilibrium conditions and includes self-consistent treatments of magnetic fields, UV irradiation, grain and PAH chemistry, and H2 excitation and cooling. Depending on the shock speed and magnetic field strength, it solves for , , or -type shocks. From each model, key outputs such as temperature, normalized upper state column density, H2 integrated intensities, shock width, and molecular abundances are provided. Comparison of JWST observational H2 data to the shock grid has been recently studied in other outflows (for e.g. see Le Gouellec , 2025; Navarro et al., 2025), demonstrating the utility of the models in constraining shock properties.
@@
The Kristensen et al. (2023) shock grid explores six key input parameters: the pre-shock density ( cm-3), the magnetic field scaling parameter (; where transverse magnetic field strength B(G) = ), the shock velocity ( km s-1, dependent on ), the external UV field strength , the H2 cosmic-ray ionization rate , and the PAH abundance ). The parameters , and are sampled in factors of 10, while and vary non-uniformly.
@@
We compared the observed extinction corrected upper state column densities from both outflows lobes (see Fig. 14) to corresponding values predicted by the models. We did not restrict the initial conditions and explored the full parameter space. For each model, the reduced was calculated to assess how well it reproduced the observed emission, and the model with minimum was initially identified as the best fit. However, we found that several models can yield values close to the minimum while having differing parameter values. To account for this degeneracy and obtain robust constraints on the likely physical conditions of the shock, we considered all models with . This value was chosen to provide a practical compromise by excluding clearly discrepant models while allowing exploration of the full parameter degeneracies in the model grid. The selection criteria resulted in 156 models that closely matched the observed emission. Fig. 10 shows the best and selected model fits on the observed data.
@@
We next explored the properties of the selected models to evaluate the most common and statistically significant parameter values. Histogram analysis of the key different parameters are shown in Fig. 10. The analysis revealed that the majority of the selected models correspond to -type shocks, indicating that these shocks are the dominant excitation mechanism in the observed regions. The magnetic field scaling factor is low, with values between 0.1 and 0.3, consistent with the absence of strong -type shock signatures. The shock velocities were found to be generally moderate, around km s-1, while the pre-shock densities range from cm-3. The external UV radiation field was distributed mostly between 10 and 100, suggesting that UV irradiation contributes alongside collisional excitation. and (PAH) parameters are not considered much as the shock modeling has a very low dependence on them (Kristensen et al., 2023).
@@
The significance of UV irradiation is highlighted in these results and suggests that, even if stellar UV is largely attenuated by the dense envelope, shock-generated UV can still be a substantial contributor to the shock structure, especially its temperature. At velocities 30-50 km s-1, post-shock temperatures reach K, leading to H2 dissociation and partial ionization. The hot post-shock gas then emits UV photons through both recombination and collisional excitation, creating a radiative precursor that enhances local photoionization. Recent study by Navarro et al. (2025) in HH 46 similarly found that -shocks accompanied by a local UV field dominate the excitation of H2. Therefore, moderate velocity dissociative -shocks can provide a natural in situ UV source while collisionally exciting both the warm and hot H2 components that dominate the observed emission. The relatively weak magnetic field further suppresses the formation of extended - type shocks. Overall, our results indicate that moderate velocity -type shocks, combined with a modest UV field, provide the best match to the observed H2 emission in L1527.

### v2

@@ 3.6 Double-peaked lines: Jet detection
Double-peaked spectral lines in protostars are linked to complex geometries and kinematics. These lines typically arise from Doppler-shifted emission caused by fast moving flows with their components moving at different velocities. High-resolution observations of many protostars have shown that forbidden emission lines often exhibit two velocity components, namely the low-velocity (LV) and HV component; see the review by Eislöffel et al. (2000). Such velocity separations can also arise in bow-shock structures, where a curved shock front produces a range of shock velocities and spatially separated emission regions (Hartigan et al., 1987). These interpretations have been supported by theoretical simulations with magnetohydrodynamic (MHD) jet models (Garcia et al., 2001; Shang et al., 2002). They have shown the presence of different velocity components through their position-velocity diagrams.
@@
Closer inspection of the L1527 MIRI/MRS spectra reveals clear double-peaked emission in three different forbidden lines: [Ar ii], [Ne iii], and [Fe ii]. The [Ne ii] line shows a very weak, unresolved profile limited to just a few pixels; it was therefore excluded. The double-peaked emission was detected only in the eastern region of the outflow. Fig. 8 shows the profiles and the morphology of the double-peaked spectra for the [Ar ii], [Ne iii], and [Fe ii] lines. In the first column, the moment 0 maps are presented with the sample aperture positions. In the middle column, we show the double-peaked spectra extracted at the sample aperture position. The spectra are color-coded in blue, green, and red for [Ar ii], [Ne iii], and [Fe ii], respectively. The line profiles clearly show two separate velocity peaks associated with the LV and HV components. A 1D Gaussian fit was performed on each velocity component to estimate their peak radial velocity, shifted to the rest frame in the LSR. In the last column of Fig. 8, we show the jet emission map associated with the HV component. The jet emission was isolated from the double-peaked line profile by extracting the HV component, thereby identifying the associated spatial region. This was achieved by integrating the line intensity over a velocity range of km s-1, with the additional criteria that the emission exhibited a Gaussian profile and a line width exceeding the spectral resolution. All three maps show an extended jet with its brightest emission closest to the protostar.
@@
The LV component observed in the double-peaked lines is distributed throughout the region and associated with the shocked gas in the general outflow or disk wind. The average radial velocity of the LV component is about km s-1, which is notably blueshifted. However, we can confirm that in regions without double-peaked profiles, the average radial velocity is redshifted and is consistent with the [Ne ii] velocity maps. Disk winds launched from a few to several au from the protostar are expected to have velocities of tens of kilometers per second (Pudritz et al., 2007; Shang et al., 2007). They are thought to shape the molecular outflow cavities by sweeping up the surrounding envelope material. Our observed LV component velocities are consistent with these predictions and support previous studies indicating that the L1527 outflow cavity is shaped by a slow disk wind entraining the ambient material (Tobin et al., 2008, 2012).
@@
The double-peaked HV component is present only in the eastern region of the outflow. The [Ar ii] line shows the strongest emission among the three detected lines. The average radial velocity of the HV component is about km s-1. The emission is neither narrow nor collimated. Rather, it is extended in width and follows the direction of the outflow axis to about 500 au. We see that the emission bends or curves toward the south in all three maps. The HV component is interpreted as evidence for the presence of a jet. Although the emission is not strictly collimated as in other class 0/I jets, its observed velocity (100 km s-1), extended scale, and detection in high-ionization lines ([Ar ii], [Ne iii]) strongly suggest shock-excited jet activity, a strong characteristic of early-stage protostellar evolution.
@@
Considering the jet is almost in the plane of sky, we can correct the radial velocity for inclination. For a value of , we obtain a true velocity of around km s-1. This value should be interpreted cautiously, as the disk is warped and the jet inclination varies significantly due to its precession and bending. Very high velocities (200 km s-1) are predicted by magneto-centrifugal launching models when jets originate very close to the protostar. X-wind models (Shu et al., 2007), which launch material near the corotation radius, naturally produce such high velocities due to the strong gravitational potential in this region. However, disk-wind models (Ferreira , 1997) can also generate comparably high velocities when launched from sub-au radii. Therefore, the high velocities observed in the jet likely indicate that its launching region lies very close to the protostar. Overall, the coexistence of both velocity components in multiple ionized lines suggests L1527 is composed of a stratified layer of gas with varying excitation conditions.
@@ 4.1 Temperature and column density maps
Physical properties such as gas temperature and column density (H2) can be inferred using excitation diagram analysis of H2 rotational transitions. The observed extinction-corrected H2 line-integrated intensities are converted into upper-state column densities normalized by its statistical weight (/). These are plotted in a logarithmic scale against their upper-state excitation energies (/). For gas in local thermal equilibrium (LTE), the gas excitation follows a Boltzmann distribution (//), and the data points in the diagram can be fit with a straight line. The gas excitation temperature is derived as the reciprocal of the slope of the line, whereas the column density is proportional to the y-axis intercept. This analysis assumes that all H2 transitions are optically thin, which is valid even for (H2)1023 cm-2 (Bitner , 2008).
@@
The MIRI/MRS spectral range covers several pure rotational H2 transitions from the S(8) to S(1) state, which are detected in most pixels in the entire field of view. This allows us to carry out pixel-by-pixel excitation diagram analysis and to obtain a complete map of , and (H2). We first smoothed the spectral cubes from different MIRI channels to the lowest common resolution, as the angular resolution varies from 02 to 10 between all the MIRI channels. We smoothed the spectral cubes to a resolution of 07, matching that of the S(1) transition, which is the lowest angular resolution in the detected H2 lines. We also regridded the cubes to the same spatial grid as the H2 S(1) data.
@@
Next, we corrected for extinction along the line of sight. The most commonly used extinction correction in MIR comes from the McClure (2009) and KP5 (Pontoppidan et al., 2024) curves, with the latter introducing slightly stronger absorption from ices and silicates. Previous analysis of other JOYS sources adopted the McClure (2009) extinction curve valid between a -band magnitude of 1 to 7 mag. However, a recent study of ices in L1527 by Slavicinska et al. (2025) showed substantial ice and silicate absorption, indicating that the KP5 curve provides a more appropriate extinction correction, which we apply in our analysis. The optical depth at 9.7m was earlier estimated to be around (Slavicinska et al., 2025). We also measured the extinction by de-reddening the S(1) to S(8) lines in the excitation diagram analysis, which yielded mag (see Fig. 13). Based on these results and previous extinction estimates from SED modeling (Tobin et al., 2008, 2010), we adopted a value of mag, which is roughly equivalent to a visual extinction of mag. We tested lower extinction values between mag and find that the intensity of the S(3) line decreases significantly and gets excluded from the fit. We also find that an H2 ortho-to-para ratio of 3 provides the best fit to our observed line intensities, assuming the H2 gas is in LTE.
@@
The excitation diagram analysis was performed using the pdrtpy222The pdrtpy project, developed by Marc Pound and Mark Wolfire, is supported by the NASA Astrophysics Data Analysis Program grant 80NSSC19K0573; the JWST-ERS program ID 1288 through grants from STScI under NASA contract NAS5-03127; and the SOFIA C+ Legacy Project through NASA award #208378 issued by USRA. python package, which is commonly employed in the analysis of JOYS sources (Gieser et al., 2023; Caratti o Garatti et al., 2024; Francis et al., 2025). Typically, the observed H2 rotational transitions trace two temperature components, often parameterized as a ”warm” and a ”hot” component in excitation diagrams. The warm component ( K) is traced by the lower excitation lines (e.g., S(0)-S(4), 4000 K) and is commonly associated with slow, non-dissociative -shocks or the cooler post-shock gas within -shocks. The hot component ( K), traced by the higher excitation transitions (e.g., S(5) and above, 5000-10000 K), arises mostly from energetic fast dissociative -shocks where collisional heating dominates. However, recent shock models (e.g., Kristensen et al., 2023) demonstrate that a single shock structure can produce both temperature components, where the hot gas represents the post-shock region and the warm gas traces the cooling layer of the same shock. Therefore, the presence of two temperature components does not necessarily indicate separate shocks but can instead reflect a thermally stratified region with distinct excitation conditions. In general, the warm component typically carries most of the H2 luminosity and therefore remains a dominant tracer of the bulk outflowing gas, while the hot component highlights localized regions of strong shock activity.
@@
At the initial step, the extinction-corrected H2 transitions at each spaxel were fit by a 1D Gaussian, and the line intensities were computed from the integral of the Gaussian fit. We set an intensity threshold of 20 MJy sr-1 to avoid any spurious detection. In Fig. B, we show example spectra for spaxels associated with the apertures in the eastern and western lobes (see Fig. 1). The spectra are plotted for different H2 transitions, and their Gaussian fits are highlighted in red and blue for the eastern and western lobes, respectively. We see that the line intensities vary across different transitions depending on their strength. In general, the lines from all transitions have bona fide profiles for the Gaussian fits without any omission.
@@
Next, the line intensities from the spectra were converted to upper-state column densities and were then plotted against their excitation temperature, as described earlier. To obtain reliable estimates from the excitation diagram, fitting was performed only if five or more transitions were detected in a spaxel. Fig. 14 shows the excitation diagram for the example spaxels in the eastern and western lobes. The data points with error bars represent the upper-state column densities of the lines in each H2 transition. They are distributed across the plot with overlays of red and blue fits for the warm and hot components, respectively. The total fit to the data points is shown as a solid green line. The estimated warm temperatures in both the spaxels are similar, with an average of about 570 K. Whereas, the hot temperature varies considerably with values of 2170 K and 3340 K for the eastern and western spaxels, respectively. The estimated warm and hot column densities in both spaxels average and .
@@
Fig. 9 presents the full temperature and column density maps for both the warm and hot components. In the temperature maps, the warm component exhibits moderate temperature variations across the outflow region, with values ranging between K. There is a noticeable gradient in the eastern lobe with temperatures increasing toward northeast. This gradient closely follows the spatial distribution of the H2 emission, particularly resembling the brighter S(5) and S(4) transitions (see Fig. 3). This suggests the presence of warm gas within the outflow cavity. The western lobe also has a gradient, though with temperatures increasing toward the southwest and with lesser resemblance to the H2 emission. In the hot component map, temperature variations are substantial, ranging between K. Both sides of the region exhibit sharp increases in temperature along the outflow structures, closely resembling the distribution of the H2 emission. These temperature variations and their spatial resemblance to the outflow morphology suggest that gas heating is strongly linked to a common excitation mechanism within the outflow.
@@
The column density maps exhibit a morphology broadly similar to the temperature maps. The warm component of the column density ranges from , whereas the hot component ranges from . The warm component is significantly denser, about two orders of magnitude higher than the hot component. In the warm component maps, regions of higher column densities show a general positive correlation with regions with higher temperatures. However, this is in contrast to the hot component maps, where regions of lower column density exhibit higher temperatures. These trends likely reflect the nature of shock heating in the outflow.
@@ 4.2 Comparison to shock models
The H2 emission from outflows in dense regions ( cm-3) is generally thought to be dominated by collisional excitation from shocks. However, UV-driven irradiation may still play a role if accretion-generated UV photons penetrate the base of the outflow cavity before being attenuated, or if fast dissociative shocks generate local UV fields that modify the shock structure and chemistry. The strong gradients in temperature, density, and ionization across shock fronts naturally produce multiple excitation components, with hotter gas populating higher-energy rotational states and warmer gas dominating lower levels. To constrain the mechanism of excitation in the outflow, we compared our observed H2 emission with recent shock models that self-consistently treat different excitation processes.
@@
We used the recent shock model grid from Kristensen et al. (2023), based on the Paris-Durham code (Flower & Pineau des Fortês , 2003; Godard et al., 2019), which computes the dynamical, thermal, and chemical structure of steady, plane-parallel shocks over 14,000 models. The code determines pre-shock equilibrium conditions and includes self-consistent treatments of magnetic fields, UV irradiation, grain and PAH chemistry, and H2 excitation and cooling. Depending on the shock speed and magnetic field strength, it solves for , , or -type shocks. From each model, key outputs such as temperature, normalized upper-state column density, H2-integrated intensities, shock width, and molecular abundances are provided. The comparison of JWST H2 observations with shock grids has recently been studied in other outflows (for e.g., see Le Gouellec , 2025; Navarro et al., 2025), demonstrating the utility of the models for constraining shock properties.
@@
The Kristensen et al. (2023) shock grid explores six key input parameters: the pre-shock density ( cm-3), the magnetic field scaling parameter (, where transverse magnetic field strength B(G) = ), the shock velocity ( km s-1, dependent on ), the external UV field strength , the H2 cosmic-ray ionization rate , and the PAH abundance ). The parameters , and are sampled in factors of 10, while and vary non-uniformly.
@@
We compared the observed extinction-corrected upper-state column densities from both outflows lobes (see Fig. 14) with the corresponding values predicted by the models. We did not restrict the initial conditions and explored the full parameter space. For each model, the reduced was calculated to assess how well it reproduced the observed emission, and the model with minimum was initially identified as the best fit. However, we found that several models can yield values close to the minimum while having differing parameter values. To account for this degeneracy and obtain robust constraints on the likely physical conditions of the shock, we considered all models with . This value was chosen to provide a practical compromise by excluding clearly discrepant models while allowing exploration of the full parameter degeneracies in the model grid. The selection criteria resulted in 156 models that closely matched the observed emission. Fig. 10 shows the best and selected model fits to the observed data.
@@
We next explored the properties of the selected models to evaluate the most common and statistically significant parameter values. Histogram analysis of the key different parameters are shown in Fig. 10. The analysis revealed that the majority of the selected models correspond to -type shocks, indicating that these shocks are the dominant excitation mechanism in the observed regions. The magnetic field scaling factor is low, with values between 0.1 and 0.3, consistent with the absence of strong -type shock signatures. The shock velocities were found to be generally moderate, around km s-1, while the pre-shock densities range from cm-3. The external UV radiation field was distributed mostly between 10 and 100, suggesting that UV irradiation contributes alongside collisional excitation. The and (PAH) parameters have minimal impact, as shock modeling does not significantly depend on them (Kristensen et al., 2023).
@@
These results highlight the significance of UV irradiation and suggest that, even if stellar UV is largely attenuated by the dense envelope, shock-generated UV remains a substantial contributor to the shock structure, especially its temperature. At velocities 30-50 km s-1, post-shock temperatures reach K, leading to H2 dissociation and partial ionization. The hot post-shock gas then emits UV photons through both recombination and collisional excitation, creating a radiative precursor that enhances local photoionization. A recent study by Navarro et al. (2025) in HH 46 similarly found that -shocks accompanied by a local UV field dominate H2 excitation. Therefore, moderate-velocity dissociative -shocks can provide a natural in situ UV source while collisionally exciting both the warm and hot H2 components that dominate the observed emission. The relatively weak magnetic field further suppresses the formation of extended - type shocks. Overall, our results indicate that moderate-velocity -type shocks, combined with a modest UV field, provide the best match to the observed H2 emission in L1527.